Consider the scalar-valued function f(x,y)=√(8y+1), and the curve C described by y=2x^2 for x∈[0,1]. The value of ∫C f(x,y)ds

is
Select one:
a. 16
b. 8
c. 19/3
d. 16/3
e. 3/32
f. 32
1
Expert's answer
2020-06-11T19:15:10-0400

Parametric equation of curve "C" is "r(t)=<t,2t^2>,0\leq t \leq1"


Here, differential length "ds" is,


"ds=\u2225r'(t) \u2225dt"


"=\u2225 <1,4t>\u2225dt"


"=\sqrt{1^2+(4t)^2}dt"


"=\sqrt{1+16t^2}dt"


Now, the line integral is evaluated as,


"\int_Cf(x,y)ds=\int_C\sqrt{8y+1}ds"


"=\int_0^{1}\sqrt{8(2t^2)+1}\sqrt{1+16t^2}dt"


"=\int_0^{1}\sqrt{1+16t^2}\sqrt{1+16t^2}dt"


"=\int_0^{1}(1+16t^2)dt"


"=[ t+16(\frac{t^3}{3})]_0^{1}"


"=1+\frac{16}{3}"

"=\frac{19}{3}"


Therefore, the line integral is "\int_Cf(x,y)ds=\frac{19}{3}".

Hence, option (c) is correct.